• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 16:03:39 on December21,2003

    Entry number 114904

    keyword=Day Shift Summary

    08:30 ep measurement

    09:30 RICH detector repair

    11:00 Start septum cooldown

    14:35 Septum cooldown completed. Vrey long cooldown Set septum current to 180.07 Amps on each side.

    14:20 Current is ramped up to desired value, and we are now ready for beam.

    14:25 Run with empty target at 5 microamps

    14:35 12C (thick) was inserted and RICH data taking started at 5 microamps